OpenStudy (anonymous):

1-3/2a This is a subtracting rational expression problem please help

OpenStudy (anonymous):

wrote it wrong *** A-3/2a

OpenStudy (anonymous):



OpenStudy (anonymous):

you can rewrite a as \[a = \frac{2a}{2a}\] right?

OpenStudy (anonymous):

im not sure im totally wack at this topic

OpenStudy (anonymous):

wait. sorry. 2a/2a is 1

OpenStudy (anonymous):

\[a = \frac{2a^2}{2a}\] that's better

OpenStudy (anonymous):

so replace the a with 2a^2/2a to get \[\frac{2a^2}{2a} - \frac 3{2a}\] do you follow?

OpenStudy (anonymous):

uhm some what

OpenStudy (anonymous):

so now you have common denominators. that means all that's left now is to subtract the numerators then copy the denominator. just like this \[\frac 32 - \frac 12 = \frac{3-1}2\] got it?

OpenStudy (anonymous):

ok so one thing I look for the lcd to get common denominator and thanks beautiful ;)